OpenStudy (anonymous):

The formula for the surface area of a cylinder is A = 2r(r + h). Find h if A = 80 and are = 5

OpenStudy (blues):

So the question gives you an equation with three variables in it. It tells you what numbers or values two of those variables are equal to. It asks you what the third variable is. So to solve something like this, you substitute the values for the A and r into the equation, then solve for h... Is that helpful?

OpenStudy (anonymous):

no

OpenStudy (blues):

OK, so the equation it gives you is: A = 2r(r + h). Is that much clear?

OpenStudy (anonymous):

i know the formulas i just need help finding the height

OpenStudy (blues):

Cool - that is exactly what we are working toward. So it gives you values for two of the variables, A = 80 and r = 5. You substitute those values for A and r into the equation, where you see A and r. Would you like to do it or would you like me to?

OpenStudy (blues):

Does that make sense?

OpenStudy (anonymous):

i'm coming out with 6, is this correct

OpenStudy (blues):

Let's see... With A and r substituted in, the formula is\[80 = (2*5)*(5 + h)\] If you do the multiplication, it is\[80 = 10(5 + h)\] and then if you expand it, it gives you\[80 = 50 + 10h\] You solve this for h by subtracting 50 from both sides, \[30 = 10 h\] And then you isolate the h by dividing both sides by 10. \[3 = h\] Is what I get.

OpenStudy (anonymous):

ok, i think i have it, thanks

OpenStudy (blues):

And now you know how to solve problems like this by yourself too?

OpenStudy (anonymous):

thanks